# Achieving Complex Image Edits via Function Aggregation with Diffusion Models

### v1

@@ Abstract
Diffusion models have demonstrated strong performance in generative tasks, making them ideal candidates for image editing. Recent studies highlight their ability to apply desired edits effectively by following textual instructions, yet two key challenges persist. First, these models struggle to apply multiple edits simultaneously, resulting in computational inefficiencies due to their reliance on sequential processing. Second, relying on textual prompts to determine the editing region can lead to unintended alterations in other parts of the image. In this work, we introduce FunEditor, an efficient diffusion model designed to learn atomic editing functions and perform complex edits by aggregating simpler functions. This approach enables complex editing tasks, such as object movement, by aggregating multiple functions and applying them simultaneously to specific areas. FunEditor is 5 to 24 times faster inference than existing methods on complex tasks like object movement. Our experiments demonstrate that FunEditor significantly outperforms recent baselines, including both inference-time optimization methods and fine-tuned models, across various metrics, such as image quality assessment (IQA) and object-background consistency.
@@ Introduction
Figure 1: (a) Results of applying basic functions, edge enhancement and object removal, on the corresponding masks. Best viewed when enlarged. (b) By applying each basic function simultaneously to different masks, we can perform complex edit functions. $\mathcal{A}$ is an aggregation function used for applying both functions in one shot.
@@
In this work, we propose FunEditor, an efficient diffusion model that performs complex edits through the aggregation of simpler functions. Under this formulation, FunEditor learns to effectively integrate multiple localized edit functions and apply them simultaneously. This is achieved by defining learnable task tokens during training and combining them during inference. For localization, each simple task requires a binary mask as input and performs cross-attention modification (Guo and Lin 2024; Simsar et al. 2023). Compared to prior editing models, FunEditor offers three unique advantages. First, during training, it only requires simple and atomic editing tasks such as object removal, harmonization, edge enhancement and etc. to learning new task tokens. During inference, by simply passing a sequence of these task tokens it applies multiple atomic functions simultaneously, requiring no energy functions. Figure 1 highlights the output of FunEditor on object movement and object resizing by aggregating object removal and edge enhancement. Second, FunEditor is data-efficient as we only require training on each atomic task individually, where training data is abundant or easy to collect. Making it the superior, low-cost option when compared to prior training-based approach like AnyDoor (Chen et al. 2024b). By training on atomic tasks it could generalize to combining them and achieve complex edits like object movement during inference. Third, FunEditor prioritizes localized editing where each simple edit function only updates a specified region in the input image. This leads to improved consistency. Furthermore, as FunEditor is compatible with pre-trained few-step diffusion models such as LCM (Luo et al. 2023), these atomic functions can be combined with pixel-space operation like copy and pasting to perform high-quality simultaneous edits in just 4 steps without the need of subject regeneration or extra optimization steps during inference as indicated in Figure 2.
To evaluate the effectiveness of FunEditor, we conducted extensive evaluations on object movement and object pasting as complex editing tasks using the COCOEE and ReS datasets. We compared FunEditor against a wide range of recently proposed editing baselines, including both inference-time optimization methods (Mou et al. 2024b) and fine-tuned models (Chen et al. 2024b). The results demonstrate that FunEditor significantly outperforms all baselines across various evaluation metrics, such as IQA and object-background consistency. Notably, FunEditor performs complex edits in just 4 steps, making it 5 to 24 times faster than the baselines in the object movement task. In addition to quantitative results, visual comparisons in Figures 5 and 6 further highlight FunEditor’s superiority in complex editing tasks.
@@
$\min_{\theta}\mathbb{E}_{y,\epsilon,t}\left[||\epsilon-\epsilon_{\theta}(z_{t},t,I_{src},T)||^{2}_{2}\right],$ (3)
@@
The COCOEE dataset is an editing benchmark compiled by Yang et al. (2022), featuring 3,500 images manually selected from the MSCOCO (Microsoft Common Objects in Context) (Lin et al. 2014) validation set. A human operator used the Segment Anything model (Kirillov et al. 2023) to extract segments from a random subset of COCOEE, identifying suitable objects and assigning diff vectors to specify their new locations, resulting in a benchmark of 100 images with corresponding masks and diff vectors for object movement tasks. In addition to our annotated COCOEE dataset, Wang et al. (2024) released the ReS dataset, comprising 100 pairs of real-world images that present challenging cases of object movement. The details of the training datasets for atomic edits can be found in the appendix.
@@
Table 1 demonstrates that FunEditor, using only $4$ steps, requires $172$ fewer NFEs and is $23$ seconds faster than the state-of-the-art method for object movement, DiffEditor (Mou et al. 2024a), on the COCOEE dataset. SelfGuidance, DragonDiffusion, and DiffEditor, rely on optimizing energy functions to guide complex edits during inference. These method requires multiple calls to the UNet, as it must perform both inpainting and blending of the object with its surrounding environment, leading to a more time-consuming process. Comparing to training-based approaches, our model is $12$ times faster than the training-based baseline AnyDoor (Chen et al. 2024b) and $5$ times faster than Paint-by-Example (Yang et al. 2022). These methods regenerate the object to fit the target position, which also demands a higher number of steps to achieve the desired result. In contrast, FunEditor achieves superior results in fewer steps by employing simpler enhancement functions, such as edge enhancement, and pixel operations, without the need for fully regenerating the object. Instead, these functions refine the composition and harmonization of the pasted object in pixel space, enabling us to achieve complex edits more efficiently. It should be noted that the studies that used synthetic data generation for object movement (Winter et al. 2024; Wang et al. 2024) did not open-source their models, therefore we could not use them as baselines to assess their latency in a multi-turn fashion.

### v2

# FunEditor: Achieving Complex Image Edits via Function Aggregation with Diffusion Models

@@ Abstract
Diffusion models have demonstrated outstanding performance in generative tasks, making them ideal candidates for image editing. Recent studies highlight their ability to apply desired edits effectively by following textual instructions, yet with two key challenges remaining. First, these models struggle to apply multiple edits simultaneously, resulting in computational inefficiencies due to their reliance on sequential processing. Second, relying on textual prompts to determine the editing region can lead to unintended alterations to the image. We introduce FunEditor, an efficient diffusion model designed to learn atomic editing functions and perform complex edits by aggregating simpler functions. This approach enables complex editing tasks, such as object movement, by aggregating multiple functions and applying them simultaneously to specific areas. Our experiments demonstrate that FunEditor significantly outperforms recent inference-time optimization methods and fine-tuned models, either quantitatively across various metrics or through visual comparisons or both, on complex tasks like object movement and object pasting. In the meantime, with only 4 steps of inference, FunEditor achieves 5–24$\times$ inference speedups over existing popular methods. The code is available at: mhmdsmdi.github.io/funeditor/.
@@ Introduction
Figure 1: (a) Results of applying the two basic functions—Edge Enhancement (center) and Object Removal (right)—using their respective masks on the input image Best viewed when enlarged. (b) Demonstration of function aggregation using the proposed method. By simultaneously applying Object Removal and Edge Enhancement on different masks, complex edits such as object shrinking (middle) and object movement (right) can be achieved. $\mathcal{A}$ represents the operation of function aggregation.
@@
In this work, we propose FunEditor, an efficient diffusion model that performs complex edits through the aggregation of simpler functions. Under this formulation, FunEditor learns to effectively integrate multiple localized edit functions and apply them simultaneously. This is achieved by defining learnable task tokens during training and combining them during inference. For localization, each simple task requires a binary mask as input and performs cross-attention modification (Guo and Lin 2024; Simsar et al. 2023). Compared to prior editing models, FunEditor offers three unique advantages. First, during training, it only requires simple and atomic editing tasks such as object removal, harmonization, edge enhancement, etc., to learn new task tokens. During inference, by passing a sequence of learned task tokens into diffusion model, FunEditor can apply multiple functions simultaneously to accomplish a complex edit, requiring no energy guidance during inference. Figure 1 highlights the output of FunEditor for object movement and object resizing by aggregating object removal and edge enhancement. Second, FunEditor is data-efficient as we only require training on each atomic task individually, for which training data is either abundant or easy to collect, making it a superior, low-cost option when compared to prior training-based approach like AnyDoor (Chen et al. 2024b). Third, FunEditor prioritizes localized editing where each simple edit function only applies to a specified region in the input image, leading to improved consistency. Lastly, as FunEditor is compatible with pre-trained few-step diffusion models such as LCM (Luo et al. 2023), these atomic functions can be combined with pixel-space operation like copy and pasting to perform high-quality simultaneous edits in only 4 steps without the need of subject regeneration or extra optimization steps during inference as indicated in Figure 2.
To evaluate the effectiveness of FunEditor, we conducted extensive evaluations on object movement and object pasting as complex editing tasks using the COCOEE and ReS datasets. We compared FunEditor against a wide range of recently proposed editing baselines, including both inference-time optimization methods (Mou et al. 2024b) and fine-tuned models (Chen et al. 2024b). The results demonstrate that FunEditor significantly outperforms all baselines across various evaluation metrics, such as IQA and object-background consistency. Notably, FunEditor performs complex edits in just 4 steps, making it 5–24$\times$ faster than the baselines in the object movement task. In addition to quantitative results, visual comparisons in Figures 5 and 6 further highlight FunEditor’s superiority in complex editing tasks.
@@
$\min_{\theta}\mathbb{E}_{y,\epsilon,t}\left[||\epsilon-\epsilon_{\theta}(z_{t},t,I_{src},I_{trg},T)||^{2}_{2}\right],$ (3)
@@
The COCOEE dataset, compiled by Yang et al. (2022), features 3,500 images manually selected from the MSCOCO (Microsoft Common Objects in Context) (Lin et al. 2014) validation set. A human operator used the Segment Anything model (Kirillov et al. 2023) to extract segments and assign diff vectors, resulting in a benchmark of 100 images with corresponding masks and diff vectors for object movement tasks. Wang et al. (2024) also released the ReS dataset, comprising 100 pairs of real-world images that present challenging cases of object movement. The details of the training datasets for atomic edits can be found in the appendix.
@@
Table 1 demonstrates that FunEditor, using only $4$ steps, requires $172$ fewer NFEs and is $23$ seconds faster than the state-of-the-art method for object movement, DiffEditor (Mou et al. 2024a), on the COCOEE dataset. SelfGuidance, DragonDiffusion, and DiffEditor, rely on optimizing energy functions to guide complex edits during inference. These method requires multiple calls to the UNet, as it must perform both inpainting and blending of the object with its surrounding environment, leading to a more time-consuming process. Comparing to training-based approaches, our model is $12$ times faster than the training-based baseline AnyDoor (Chen et al. 2024b) and $5$ times faster than Paint-by-Example (Yang et al. 2022). These methods regenerate the object to fit the target position demanding a higher number of steps to achieve the desired result. In contrast, FunEditor achieves superior results in fewer steps by employing simpler enhancement functions, such as edge enhancement, and pixel operations, without the need for fully regenerating the object. Instead, these functions refine the composition and harmonization of the pasted object in pixel space, enabling us to achieve complex edits more efficiently. It should be noted that the studies that used synthetic data generation for object movement (Winter et al. 2024; Wang et al. 2024) did not open-source their models, therefore we could not use them as baselines to assess their latency in a multi-turn fashion.